OpenStudy (anonymous):

? < square root of 20 < ?

OpenStudy (saifoo.khan):

\[\Huge \sqrt{20}\]

OpenStudy (anonymous):

\[? < \sqrt{20}< ?\]

OpenStudy (saifoo.khan):

\[1 < \sqrt{20}<\sqrt{30}\]

OpenStudy (anonymous):

\[2\sqrt{5}\]

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

It said thats wrong

OpenStudy (saifoo.khan):

what are the options?

OpenStudy (anonymous):

there is none...it just said what number belongs in the first ?

OpenStudy (anonymous):

try 1 and 3

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

answer is \[4<\sqrt{20}<5\]

OpenStudy (anonymous):

YAY!!!!! YOUR RIGHT!!!! (((:

OpenStudy (saifoo.khan):

Sat rules OS.

OpenStudy (saifoo.khan):

HAIL SATELLITE/

OpenStudy (anonymous):

lmao im so dum

OpenStudy (anonymous):

Lol Thank youu!!!!!